# SQLong: Enhanced NL2SQL for Longer Contexts with LLMs

### v1

@@
Abstract.
Open-weight large language models (LLMs) have significantly advanced performance in the Natural Language to SQL (NL2SQL) task. However, their effectiveness diminishes when dealing with large database schemas, as the context length increases. To address this limitation, we present SQLong, a novel and efficient data augmentation framework designed to enhance LLM performance in long-context scenarios for the NL2SQL task. SQLong generates augmented datasets by extending existing database schemas with additional synthetic CREATE TABLE commands and corresponding data rows, sampled from diverse schemas in the training data. This approach effectively simulates long-context scenarios during finetuning and evaluation. Through experiments on the Spider and BIRD datasets, we demonstrate that LLMs finetuned with SQLong-augmented data [^1] The data will be made publicly available upon paper acceptance. significantly outperform those trained on standard datasets. These imply SQLong’s practical implementation and its impact on improving NL2SQL capabilities in real-world settings with complex database schemas.
LLMs, NL2SQL, Long Contexts
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Conference acronym ’25; 2025; Woodstock, NYccs: Computing methodologies Natural language processingccs: Computing methodologies Neural networksccs: Information systems Social networks
1. Introduction
Figure 1. Our proposed SQLong Pipeline.
The NL2SQL task focuses on translating natural language questions into SQL queries, enabling non-experts to interact with databases seamlessly (Deng et al., 2022). Recent advances leverage LLMs, finetuned on structured input prompts (e.g., task instructions, database schema, and natural language question), to achieve state-of-the-art performance on benchmarks like Spider (Yu et al., 2018) and BIRD (Li et al., 2023) (Yang et al., 2024a; Liu et al., 2024).
Despite significant progress, a critical challenge persists: LLMs finetuned on existing benchmarks still struggle with large database schemas due to limited context handling. Current datasets primarily feature small schemas, failing to represent real-world complexities. Additionally, the absence of publicly available large-schema datasets further hinders progress. Addressing this, we propose SQLong, a data augmentation framework designed to enhance LLM performance in long-context NL2SQL tasks by extending schemas to meet predefined context thresholds.
@@ 1. Introduction
In this paper, we focus on demonstrating that SQLong-augmented models outperform their unaugmented counterparts across varying context lengths. While direct comparisons to retrieval-augmented generation (RAG) schema linking are beyond this paper’s scope, our findings suggest combining SQLong with RAG could unlock further gains.
Our main contributions include:
  • Introducing long-context NL2SQL: A challenging new task for evaluating LLM performance on large database schemas.
  • SQLong pipeline: A novel, scalable data augmentation approach for generating long-context training and test datasets.
  • Empirical insights: Comprehensive experiments validating SQLong’s effectiveness in enhancing LLM robustness and accuracy in long-context scenarios.
  • Resource sharing: Plans to release SQLong datasets and code to support further research.
2. The Proposed SQLong Pipeline
Figure 2. Prompt template for the NL2SQL task.
@@ 2. The Proposed SQLong Pipeline
The SQLong pipeline comprises three main steps:
  1. (1)
@@
    Schema Collection. We collect all $\mathsf{CREATE\ TABLE}$ commands and three sample data rows for each table from the training database schemas, compiling them into a comprehensive schema set.
  2. (2)
@@
    Schema Augmentation. For each training pair, consisting of an input prompt (task instructions, database schema, natural language question) and its target SQL query, SQLong randomly samples items from the schema set. These sampled items contain table names distinct from those in the given database schema. The sampled items are combined with the original schema, and the resulting schema is randomly shuffled to produce a new, long-context database schema. This shuffling introduces variability in the positions of the original tables and columns.
  3. (3)
@@
    Long-Context Prompt Generation. SQLong generates an augmented input prompt in the format of task instructions, the long-context database schema, and the natural language question, while keeping the target SQL query unchanged. It ensures that the combined length of the long-context input prompt and the target SQL query does not exceed a predefined context length (e.g., 32k tokens), maintaining compatibility with the model’s tokenizer constraints.
@@
By systematically extending and diversifying the context, SQLong enhances the robustness and effectiveness of LLMs in handling long-context NL2SQL tasks. We summarise the steps involved in SQLong in Algorithm 1.
@@ 26      $augmented\_long\_context\_schema\leftarrow$ shuffle_list($augmented\_schema\ \cup database\_schema$)
3. Evaluation
We assess the effectiveness of our proposed SQLong model in enhancing NL2SQL performance across both short-context and long-context scenarios.
3.1. Experimental Setup
Figure 3. Statistics of input prompt lengths with respect to Llama-3.1-8B-Instruct’s tokenizer (left) and CodeQwen1.5-7B-Chat’s tokenizer (right) on the original BIRD-dev set. Similarly, the maximum input prompt lengths for the original Spider-related sets are approximately 2,000 tokens for Llama-3.1-8B-Instruct’s tokenizer and 2,500 tokens for CodeQwen1.5-7B-Chat’s tokenizer.
Dataset #DB #tables #training #dev #test
Spider 200 5 $\pm$ 3 6,712 1,034 2,019
Spider-syn 200 5 $\pm$ 3 6,712 1,034
Spider-realistic 200 5 $\pm$ 3 6,712 508
BIRD 98 7 $\pm$ 3 9,428 1,534
Table 1. Statistics of the experimental datasets. #DB denotes the number of databases. #tables denotes the mean and standard deviation of numbers of tables in the databases.
Qwen2-72B-Instruct 82.7 80.7 73.0 82.9 53.7 74.6
CodeQwen1.5-7B-Chat 76.4 70.1 62.7 75.1 44.3 65.7
[…]
Finetuned with SQLong 83.4 79.7 71.2 81.3 53.3 73.8
Llama-3.1-70B-Instruct 80.7 78.0 73.0 83.7 61.5 75.4
Llama-3.1-8B-Instruct 71.1 63.8 61.0 65.7 40.9 60.5
[…]
Finetuned with SQLong 83.2 78.0 73.1 81.8 53.3 73.9
Table 2. Execution-match accuracy results (in %) across different datasets and model configurations. Finetuning with SQLong consistently improves performance, with the best results highlighted in bold.
Length CodeQwen1.5-7B-Chat Llama-3.1-8B-Instruct
Spider-related BIRD-dev Spider-related BIRD-dev
8k 37 $\pm$ 4 35 $\pm$ 8 48 $\pm$ 5 48 $\pm$ 8
16k 72 $\pm$ 6 76 $\pm$ 8 94 $\pm$ 7 102 $\pm$ 9
24k 107 $\pm$ 7 118 $\pm$ 8 141 $\pm$ 8 157 $\pm$ 9
32k 142 $\pm$ 8 159 $\pm$ 9 186 $\pm$ 8 211 $\pm$ 9
40k 177 $\pm$ 8 200 $\pm$ 9 233 $\pm$ 9 269 $\pm$ 9
48k 212 $\pm$ 9 242 $\pm$ 9 279 $\pm$ 9 320 $\pm$ 10
56k 247 $\pm$ 9 283 $\pm$ 9 326 $\pm$ 9 374 $\pm$ 9
64k 283 $\pm$ 9 324 $\pm$ 9 372 $\pm$ 8 429 $\pm$ 9
128k 551 $\pm$ 4 639 $\pm$ 7 725 $\pm$ 9 843 $\pm$ 8
Table 3. Mean and standard deviation statistics of the numbers of tables in input prompts for our augmented long-context test sets with respect to each model’s tokenizer.
Datasets
Importantly, the long-context test sets are constructed with distinct database schema alignments. To build Spider-based long-context test sets, we use the database schemas from the BIRD training set, whereas for the BIRD-dev long-context test sets, we use the database schemas from the Spider training set. This ensures a robust evaluation across diverse schema configurations and context lengths. The data statistics of the experimental datasets are presented in Figure 3 and Tables 1 and 3.
Baseline Models and Evaluation Metrics
Training Protocol
Inference Protocol
Figure 4. Execution-match accuracy (in %) for Llama-3.1 (top) and Qwen (bottom) families on Spider-test (left) and BIRD-dev (right) long-context test sets.
3.2. Main Results
Performance on Original Datasets
Table 2 summarizes the results on the original development and test sets, comparing base models with larger LLMs such as Llama-3.1-70B-Instruct (Dubey et al., 2024) and Qwen2-72B-Instruct (Yang et al., 2024b). Models finetuned using long-context augmentation via SQLong consistently outperform their counterparts finetuned on original contexts. On average, SQLong delivers an absolute improvement of over 2.2% across five benchmark datasets. Additionally, SQLong-finetuned models achieve performance comparable to much larger LLMs on specific datasets, showcasing the scalability and efficiency of the approach.
Performance on Long-Context Datasets
Figure 4 illustrates the experimental results on long-context test sets. Across all datasets, models finetuned with SQLong demonstrate superior performance compared to those trained without SQLong. For instance, on the Spider-test datasets with 8k and 24k context lengths, the Llama-3.1-8B-Instruct model achieves outstanding results of 77.1% and 72.3%, reflecting absolute gains of 7.2% and 13.3%, respectively. Notably, the SQLong-finetuned Llama-8B model outperforms the larger Llama-70B model on 41 out of 45 long-context test sets, with minor exceptions on Spider-realistic 8k and BIRD-dev 8k, 16k, and 24k sets. Similar performance trends are observed with the Qwen models.
Figure 5. Robust impact of fine-tuned models.
Positional robustness
To evaluate the positional robustness of fine-tuned models, we conduct an experiment where each original database schema is placed at varying positions within the input prompt to assess the models’ ability to detect them.
4. Conclusion and Future Work

### v2

@@
Abstract
Open-weight large language models (LLMs) have significantly advanced performance in the Natural Language to SQL (NL2SQL) task. However, their effectiveness diminishes when dealing with large database schemas, as the context length increases. To address this limitation, we present SQLong, a novel and efficient data augmentation framework designed to enhance LLM performance in long-context scenarios for the NL2SQL task. SQLong generates augmented datasets by extending existing database schemas with additional synthetic CREATE TABLE commands and corresponding data rows, sampled from diverse schemas in the training data. This approach effectively simulates long-context scenarios during finetuning and evaluation. Through experiments on the Spider and BIRD datasets, we demonstrate that LLMs finetuned with SQLong-augmented data significantly outperform those trained on standard datasets. These imply SQLong’s practical implementation and its impact on improving NL2SQL capabilities in real-world settings with complex database schemas. [^1] Table Representation Learning Workshop at ACL 2025
SQLong: Enhanced NL2SQL for Longer Contexts with LLMs
Dai Quoc Nguyen, Cong Duy Vu Hoang, Duy Vu, Gioacchino Tangari Thanh Tien Vu, Don Dharmasiri, Yuan-Fang Li, Long Duong Oracle Corporation {dai.nguyen,vu.hoang,duy.vu,gioacchino.tangari}@oracle.com {thanh.v.vu,don.dharmasiri,yuanfang.li,long.duong}@oracle.com
1 Introduction
The NL2SQL task focuses on translating natural language questions into SQL queries, enabling non-experts to interact with databases seamlessly (Deng et al., 2022). Recent advances leverage LLMs, finetuned on structured input prompts (e.g., task instructions, database schema, and natural language question), to achieve state-of-the-art performance (Yang et al., 2024b; Liu et al., 2024) on benchmarks such as Spider (Yu et al., 2018) and BIRD (Li et al., 2023). Despite significant progress, a critical challenge persists: LLMs finetuned on existing benchmarks still struggle with large database schemas due to limited context handling. Current datasets primarily feature small schemas, failing to represent real-world complexities. Additionally, the absence of publicly available large-schema datasets further hinders progress. Addressing this, we propose SQLong, a data augmentation framework designed to enhance LLM performance in long-context NL2SQL tasks by extending schemas to meet predefined context thresholds.
Figure 1: Our proposed SQLong Pipeline.
@@ 1 Introduction
In this paper, we focus on demonstrating that SQLong-augmented models outperform their unaugmented counterparts across varying context lengths. While direct comparisons to retrieval-augmented generation (RAG) schema linking are beyond this paper’s scope, our findings suggest combining SQLong with RAG could unlock further gains. Our main contributions include:
$\bullet$ Introducing long-context NL2SQL: A challenging new task for evaluating LLM performance on large database schemas.
$\bullet$ SQLong pipeline: A novel, scalable data augmentation approach for generating long-context training and test datasets.
$\bullet$ Empirical insights: Comprehensive experiments validating SQLong’s effectiveness in enhancing LLM robustness and accuracy in long-context scenarios.
$\bullet$ Resource sharing: Plans to release SQLong datasets and code to support further research.
2 The Proposed SQLong Pipeline
Figure 2: Prompt template for the NL2SQL task.
@@ 2 The Proposed SQLong Pipeline
The SQLong pipeline has three main steps:
@@
1. Schema Collection. We collect all $\mathsf{CREATE\ TABLE}$ commands and three sample data rows for each table from the training database schemas, compiling them into a comprehensive schema set.
@@
2. Schema Augmentation. For each training pair, consisting of an input prompt (task instructions, database schema, natural language question) and its target SQL query, SQLong randomly samples items from the schema set. These sampled items contain table names distinct from those in the given database schema. The sampled items are combined with the original schema, and the resulting schema is randomly shuffled to produce a new, long-context database schema. This shuffling introduces variability in the positions of the original tables and columns.
@@
3. Long-Context Prompt Generation. SQLong generates an augmented input prompt in the format of task instructions, the long-context database schema, and the natural language question, while keeping the target SQL query unchanged. It ensures that the combined length of the long-context input prompt and the target SQL query does not exceed a predefined context length (e.g., 32k tokens), maintaining compatibility with the model’s tokenizer constraints.
@@
By systematically extending and diversifying the context, SQLong enhances the robustness and effectiveness of LLMs in handling long-context NL2SQL tasks. We summarise the steps involved in SQLong in Algorithm 1 in Appendix A.1.
3 Evaluation
We assess the effectiveness of our proposed SQLong model in enhancing NL2SQL performance in both short-context and long-context scenarios.
3.1 Experimental Setup
Datasets
Importantly, the long-context test sets are constructed with distinct database schema alignments. To build Spider-based long-context test sets, we use the database schemas from the BIRD training set, whereas for the BIRD-dev long-context test sets, we use the database schemas from the Spider training set. This ensures a robust evaluation across diverse schema configurations and context lengths. The data statistics of the experimental datasets are presented in Figure 3 and Tables 1 and 2.
Figure 3: Statistics of input prompt lengths with respect to Llama-3.1-8B-Instruct’s tokenizer (left) and CodeQwen1.5-7B-Chat’s tokenizer (right) on the original BIRD-dev set. Similarly, the maximum input prompt lengths for the original Spider-related sets are approximately 2,000 tokens for Llama-3.1-8B-Instruct’s tokenizer and 2,500 tokens for CodeQwen1.5-7B-Chat’s tokenizer.
Dataset #DB #tables #training #dev #test
Spider 200 5 $\pm$ 3 6,712 1,034 2,019
Spider-syn 200 5 $\pm$ 3 6,712 1,034
Spider-realistic 200 5 $\pm$ 3 6,712 508
BIRD 98 7 $\pm$ 3 9,428 1,534
Table 1: Statistics of the experimental datasets. #DB denotes the number of databases. #tables denotes the mean and standard deviation of numbers of tables in the databases.
Length CodeQwen1.5-7B-Chat Llama-3.1-8B-Instruct
Spider-related BIRD-dev Spider-related BIRD-dev
8k 37 $\pm$ 4 35 $\pm$ 8 48 $\pm$ 5 48 $\pm$ 8
16k 72 $\pm$ 6 76 $\pm$ 8 94 $\pm$ 7 102 $\pm$ 9
24k 107 $\pm$ 7 118 $\pm$ 8 141 $\pm$ 8 157 $\pm$ 9
32k 142 $\pm$ 8 159 $\pm$ 9 186 $\pm$ 8 211 $\pm$ 9
40k 177 $\pm$ 8 200 $\pm$ 9 233 $\pm$ 9 269 $\pm$ 9
48k 212 $\pm$ 9 242 $\pm$ 9 279 $\pm$ 9 320 $\pm$ 10
56k 247 $\pm$ 9 283 $\pm$ 9 326 $\pm$ 9 374 $\pm$ 9
64k 283 $\pm$ 9 324 $\pm$ 9 372 $\pm$ 8 429 $\pm$ 9
128k 551 $\pm$ 4 639 $\pm$ 7 725 $\pm$ 9 843 $\pm$ 8
Table 2: Mean and standard deviation statistics of the numbers of tables in input prompts for our augmented long-context test sets with respect to each model’s tokenizer.
Baseline Models and Evaluation Metrics
Qwen2-72B-Instruct 82.7 80.7 73.0 82.9 53.7 74.6
CodeQwen1.5-7B-Chat 76.4 70.1 62.7 75.1 44.3 65.7
[…]
   Finetuned with SQLong 83.4 79.7 71.2 81.3 53.3 73.8
Llama-3.1-70B-Instruct 80.7 78.0 73.0 83.7 61.5 75.4
Llama-3.1-8B-Instruct 71.1 63.8 61.0 65.7 40.9 60.5
[…]
   Finetuned with SQLong 83.2 78.0 73.1 81.8 53.3 73.9
Table 3: Execution-match accuracy results (in %) across different datasets and model configurations. Finetuning with SQLong consistently improves performance, with the best results highlighted in bold.
Training Protocol
Inference Protocol
Figure 4: Execution-match accuracy (in %) for Llama-3.1 (top) and Qwen (bottom) families on Spider-test (left) and BIRD-dev (right) long-context test sets.
3.2 Main Results
Performance on Original Datasets
Table 3 summarizes the results on the original development and test sets, comparing base models with larger LLMs such as Llama-3.1-70B-Instruct (Dubey et al., 2024) and Qwen2-72B-Instruct (Yang et al., 2024a). Models finetuned using long-context augmentation via SQLong consistently outperform their counterparts finetuned on original contexts. On average, SQLong delivers an absolute improvement of over 2.2% across five benchmark datasets. Additionally, SQLong-finetuned models achieve performance comparable to much larger LLMs on specific datasets, showcasing the scalability and efficiency of the approach.
Performance on Long-Context Datasets
Figure 4 illustrates the experimental results on long-context test sets. The full details are presented in Tables 4 and 5 in Appendix A.2. Across all datasets, models finetuned with SQLong demonstrate superior performance compared to those trained without SQLong. For instance, on the Spider-test datasets with 8k and 24k context lengths, the Llama-3.1-8B-Instruct model achieves outstanding results of 77.1% and 72.3%, reflecting absolute gains of 7.2% and 13.3%, respectively. Notably, the SQLong-finetuned Llama-8B model outperforms the larger Llama-70B model on 41 out of 45 long-context test sets, with minor exceptions on Spider-realistic 8k and BIRD-dev 8k, 16k, and 24k sets. Similar performance trends are observed with the Qwen models.
Figure 5: Robust impact of fine-tuned models.
Positional robustness
We conduct an experiment wherein each original database schema is placed at different positions within the input prompt, assessing the models’ ability to detect it regardless of its location.
4 Conclusion and Future Work
Appendix A Appendix
A.1 The algorithm steps in SQLong
@@ 26      $augmented\_long\_context\_schema\leftarrow$ shuffle_list($augmented\_schema\ \cup database\_schema$)
A.2 Full execution-match accuracy results for all long-context test sets
Model Context Dataset Average
length Spider-dev Spider-realistic Spider-syn Spider-test BIRD-dev across 45 sets
Llama-3.1-8B-Instruct 8k 61.9 53.5 45.1 61.7 35.1
16k 58.5 47.0 38.9 54.7 31.3
24k 53.2 43.1 32.7 51.2 28.1
32k 49.6 42.9 29.9 47.0 26.2
40k 48.7 38.4 28.4 43.2 25.2 37.2
48k 46.9 35.8 24.9 43.8 23.9
56k 45.5 32.1 23.8 41.5 23.6
64k 42.6 33.1 22.5 39.0 21.6
128k 28.0 17.9 10.3 26.1 15.7
Our model fine-tuned 8k 71.7 63.4 49.3 69.9 43.2
       Without SQLong 16k 66.6 54.7 39.9 62.9 37.9
24k 63.6 52.4 35.5 59.0 35.2
32k 59.4 48.0 33.1 56.2 32.6
40k 57.0 45.1 30.2 52.4 33.5 43.8
48k 55.9 43.7 28.0 51.1 30.5
56k 52.5 40.4 25.7 48.3 28.8
64k 51.4 40.9 25.3 45.7 29.3
128k 34.7 23.6 13.5 30.6 18.8
Our model fine-tuned 8k 77.4 67.1 61.7 77.1 49.3
       With SQLong 16k 75.2 66.1 53.4 74.1 46.0
24k 71.8 64.2 50.0 72.3 43.8
32k 68.3 61.6 46.5 68.6 43.9
40k 67.5 62.8 44.9 66.8 42.7 54.8
48k 66.9 56.7 40.2 65.5 41.2
56k 63.3 52.6 38.4 63.5 39.7
64k 61.3 52.2 39.3 62.1 37.6
128k 43.0 33.7 21.7 41.6 24.5
Llama-3.1-70B-Instruct 8k 73.9 67.3 55.0 76.9 52.3
16k 67.7 59.4 48.9 70.2 47.1
24k 62.4 54.9 43.8 68.4 44.2
32k 60.9 49.6 41.7 63.7 40.8
40k 59.0 52.6 37.4 60.8 39.5 48.5
48k 57.6 46.9 35.0 58.3 37.1
56k 55.3 46.3 32.3 57.8 36.5
64k 55.0 43.9 31.7 54.1 34.4
128k 28.0 25.6 12.3 23.4 11.0
Table 4: Execution-match accuracy results (in %) on the augmented long-context test sets with respect to the Llama-3.1 model family.
Model Context Dataset Average
length Spider-dev Spider-realistic Spider-syn Spider-test BIRD-dev across 45 sets
CodeQwen1.5-7B-Chat 8k 61.7 49.6 38.1 60.6 32.4
16k 55.9 42.1 30.7 51.7 29.3
24k 51.5 37.8 27.9 47.9 25.6
32k 48.0 30.9 22.8 45.0 23.7
40k 46.7 28.9 21.0 40.6 21.1 31.7
48k 42.4 27.8 18.7 39.6 19.6
56k 36.4 24.0 17.5 34.0 17.3
64k 36.4 21.3 15.8 32.7 16.5
128k 19.2 7.9 6.4 16.1 7.1
Our model fine-tuned 8k 68.9 57.1 39.5 66.6 39.0
       Without SQLong 16k 62.6 51.4 31.8 57.8 35.7
24k 57.6 49.0 29.3 55.1 30.1
32k 53.0 41.5 25.6 51.3 29.6
40k 53.7 38.4 23.5 49.3 26.0 37.8
48k 48.7 34.6 22.3 45.3 24.4
56k 44.5 33.1 20.9 42.3 23.3
64k 43.8 30.3 18.4 39.8 22.1
128k 26.1 15.6 9.2 24.1 9.4
Our model fine-tuned 8k 75.9 65.7 53.2 73.4 46.9
       With SQLong 16k 72.9 62.6 46.6 68.6 42.5
24k 68.9 58.5 43.0 64.2 40.3
32k 67.5 54.3 40.0 62.9 39.2
40k 63.4 53.7 37.4 59.1 37.0 50.2
48k 63.9 52.8 35.3 58.0 33.4
56k 60.3 51.0 33.6 56.1 32.0
64k 60.6 52.4 31.0 54.1 30.8
128k 43.4 33.7 19.4 40.1 19.4
Qwen2-72B-Instruct 8k 70.6 63.4 47.2 70.9 45.9
16k 69.1 58.7 40.6 65.5 40.2
24k 60.9 53.3 34.1 56.5 37.1
32k 59.6 45.5 31.1 53.9 35.8
40k 55.8 45.7 29.5 52.7 35.4 44.2
48k 52.3 43.7 27.8 49.2 32.8
56k 50.8 39.4 27.6 45.6 32.5
64k 47.3 34.6 25.1 45.0 33.0
128k 36.8 28.3 18.6 33.3 24.9
Table 5: Execution-match accuracy results (in %) on the augmented long-context test sets with respect to the Qwen mdoel family.